X is a discrete random variable that takes values (1,2,3,4,5). If P(X=3)=0.5 and F(4)=0.85. What is the P(X=5)?
Answer #1

Solution :

Given that,

P(X=3)=0.5 and F(4)=0.85

So,

P(X = 5) = 1 - F(4) = 1 - 0.85 = 0.15